OpenStudy (anonymous):

Simplify the expressions below. Write the final product in standard form and show your work to receive full credit. 2x4(4x2 + 3x + 1) (4x – 3)(2x2 – 7x + 1) (x2 + 4x – 3)(2x2 + x + 6)

OpenStudy (anonymous):

Standard form meaning... simplified?

OpenStudy (anonymous):

Standard form means like when the exponents are listed from high to low in order in an equation

OpenStudy (anonymous):

Ohhh, then wouldn't you multiply out first and then order it? 2x^4(4x^2+3x+1) 8x^6+6x^5+2x^4 (4x – 3)(2x^2 – 7x + 1) 8x^3-28x^2+4x-6x^2+21x-3 8x^3-34x^2+25x-3 (x^2 + 4x – 3)(2x^2 + x + 6) 2x^4+x^3+6x^2+8x^3+4x^2+24x-6x^2-3x-18 2x^4+9x^3+4x^2+21x-18

OpenStudy (anonymous):

Feel free to check my work. It's late here and all the numbers just start to blend together after a while xD

OpenStudy (anonymous):

2x^4=(4x^2+3x+1)=8x^6+6x^5+2x^4 (4x – 3)(2x2 – 7x + 1)=8x^3-34x^2+25x-3 (x2 + 4x – 3)(2x2 + x + 6)=2x^4+9x^3+4x^2+21x-18